OpenStudy (anonymous):

(D-1)x + Dy = 0 (D+1)x + (2D+2)y = 0 How can i solve these equ. in MATLAB ?

OpenStudy (anonymous):

by matrices

OpenStudy (anonymous):

um, i'm only in high school, so i haven't read much about this type of equations, but from what i know, here's how i solved it, even if it's probably false: 1. let's start with : (D-1)x+ Dy = 0 according to it, we find D= (x)/(x+y) 2. now let's replace D in the second equation, it becomes like this: (D+1)x+ (2D+2)y= 0 Dx + x + 2Dy + 2y= 0 D(x+2y) + x + 2y= 0 (x+ 2y) (D+1)= 0 (now we replace): (x+ 2y) ([x/(x+y)] + 1)= 0 (x+2y) ( [x(x+y)]/[x+y] )= 0 [x(x+y)(x+2y)]/(x+y)= 0 (x²+xy)(x+2y) =0 x^3 + 2x²y+ x²y+ 2xy² =0 x²+y(3x+2y) =0 x²+ 3xy+ 2y² =0 x²+ y²+ 2xy +xy + y² =0 (x+y)²+ y(x+y)=0 therefor: x=0 y=0 I don't know if 'im correst or not, i'm only 15 and i live in Africa!!